1/2

Christian Morgenstern,  Christian Morgenstern,  Christian Morgenstern, Christian Morgenstern,
Le loup-garou  Varulven  El Hechicero  The Werewolf
translated by R. Platteau  translated by Sven Collberg  translated by Hans Eiseneck translated by Alexander Gross

Un loup-garou, une certaine nuit, Laissa sa moitié et puis se rendit Auprès de la tombe d’un maître d’école Et lui dit, “ sur moi dis tout, c’est ton rôle ! ”

En varulv lopp från fru m.m. en vacker afton samt begav sig till en bymagisters grav och bad: “Var snäll och konju-

gera.

Un hechicero, con esposa vino un día a la fosa de su maestro, y al profesor pidió: “¿Me cuentas? ¡Por

favor!”

A Werewolf, troubled by his name, Left wife and brood one night and came To a hidden graveyard to enlist The aid of a long-dead philologist.

Un loup-garou, une certaine nuit, Laissa sa moitié et puis se rendit Auprès de la tombe d’un maître d’école Et lui dit, “ sur moi dis tout, c’est ton rôle ! ” mig!” Byskolläraren stod upp på blecknamnsskyltens

mässingknopp med svar till den, som satt på

pass med tåligt korslagd tass på

tass.

El profesor sin vacilar subió y comenzó a hablar. Desde allá de su letrero se dirigió al hechicero:

“Oh sage, wake up, please don’t berate

me,” He howled sadly, “Just conjugate me.” The seer arose a bit unsteady Yawned twice, wheezed once, and then

was ready.

“ - On peut à ton sujet puisque nous y sommes Se demander encor bien des choses en somme. Loup-garoù ? mais avant Il faut se dire : loup-garquand ?

“Jag var-ulv 1:a pers. i sing., du, han, hon, den, det var-ulv.

Så vi voro-ulv, I voren. På de voro-ulv kom ingenting.”

“El hechiuno” le explica, “el hechidos” después indi-

ca, “el hechitres” dice al fin, “con eso basta ¡chiquitín!”

“Well, ‘Werewolf’ is your plural past, While ‘Waswolf’ is singularly cast: There’s ‘Amwolf’ too, the present tense, And ‘Iswolf,’ ‘Arewolf’ in this same sen-

se.”

Loup-garcomment ? et puis Bien évidemment loup-garqui ? ” Toutes ces questions semblèrent au loup bien

absconses, Car en fait ce qu’il voulait c’était les réponses.

Varulven myste gladåt hur pers. hade följts av sing. och

plur. “Fast”, bad han, “kan det inte

ges ens en enda pers. i tempus pre-

sens?”

El hechicero, conmovido por lo que capta su oído, al maestro ruega, encanta-

do, que le adjunte un quebrado.

“I know that–I’m no mental cripple– The future form and participle Are what I crave,” the beast replied. The scholar paused–again he tried:

…  …  …  …

2/2

“ - Point de réponses, dit le puits de savoir, Etre imaginaire, nul ne peut te voir ! ” A ces mots l’animal éclata tout en larmes. “ - Oh!, même les baisers de ma douce moitié Bekänna måste vår magister, att därvidlag fanns vissa bri-

ster. En varulv sine imperfecto var contradictio in adjecto.

Pero el profesor confiesa que nada sabe con certeza: “Hay muchos hechos. Hechi-

ceros únicamente hay enteros.”

“A ‘Will-be-wolf?’ It’s just too long: ‘Shall-be-wolf?’ ‘Has-been-wolf?’ Utterly

wrong! Such words are wounds beyond all

suture– I’m sorry, but you have no future.”

Ne pourrons plus me faire me sentir tout en-

tier. Rester un loup-garou est un sort qui m’alarme.

” Mais comme le loup était bien poli Il remercia et puis repartit.

Varulven satt med tårad lins. Han visste att han fanns och

finns. Därtill kom också fru m.m. som även måste existera.

El brujo se volvió muy tri-

ste: “Mi cara mitad, ¿no la vi-

ste?” Ya que no era sabio con gratitud se despidió.

The Werewolf knew better–his sons still

slept At home, and homewards now he crept, Happy, humble, without apology For such folly of philology.

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into French by R. Platteau.

Arg-825-1655 (2013-02-13 21:56:38)

Tiu ĉi francigo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into Swedish by Sven Collberg.

Arg-825-1654 (2013-02-13 17:00:43)

Tiu ĉi francigo troviĝas ne la reterjo http: // bertilow. com/ literaturo/ lupfantomo. html# sv .

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into hispana by Hans Eiseneck.

Arg-825-1656 (2013-02-13 18:46:00)

Tiu ĉi trsduko en la kastilian lingvon troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .

Translation of the German poem “Der Werwolf” by Christian Morgenstern (Kristiano Matenstelo, ∗1871-05-06 – †1914-03-31) into English by Alexander Gross.

Arg-825-1657 (2013-02-13 18:39:45)

Tiu ĉi angligo troviĝas en http: // bertilow. com/ literaturo/ lupfantomo. html .